# Event-Aided Sharp Radiance Field Reconstruction for Fast-Flying Drones

### v1

@@ III-A Trajectory Representation and Pose Refinement
where $\gamma_{L_{t}}(\cdot)$ denotes a frequency-encoded representation of time, following the sinusoidal positional encoding introduced in NeRF [29]. The network $f_{\phi}$ maps this embedding to the residual pose $\Delta\mathbf{T}_{\text{refine}}(t)$, computed from the predicted translation $\mathsf{v}(t)\in\mathbb{R}^{3}$ and quaternion $\mathsf{q}(t)\in\mathbb{R}^{4}$ vectors.
@@ III-B Image Supervision via Motion Blur Rendering
Given a pixel $\mathbf{u}$ in a frame captured at time $t_{i}$, where $t_{i}$ denotes the center of the exposure interval for image $i$, we define as $\mathbf{r}(\mathbf{u},t_{i})$ the ray that originates from the camera center of the refined camera pose $\mathbf{T}_{\text{ref}}(t_{i})$ and passes through the pixel $\mathbf{u}$. We estimate the color of the blurred pixel $\mathbf{u}$ observed in $\mathbf{C}^{\text{blur}}_{i}$ by averaging the colors the set of rays $\mathbf{r}(\mathbf{u},t_{i})$ observe as the camera moves during the exposure interval $t_{i}-\tau/2,\ t_{i}+\tau/2]$. This simulates the physical image formation process, where the observed pixel color corresponds to the integral of the latent sharp radiance accumulated over the exposure duration [27].
@@
where $w_{m}$ denotes the blending weight for the $m$-th timestamp and satisfies $\sum_{m=1}^{M}w_{m}=1$, and $g(\cdot)$ a gamma correction function. Similar to [37], we use uniform weights to blend the rendered colors over time.
Shortening the rendered pixel in Eq. 5 as $\hat{\mathbf{C}}^{\text{blur}}_{\mathbf{r}}$, we finally compare the synthetic and observed blurry pixels over a batch $\mathcal{B}_{b}$ of pixel-timestamp pairs $(\mathbf{u}_{k},t_{k})$:
@@ IV-A Implementation Details
Training.  We implement our code using PyTorch. We train using a batch size of 1024 rays for the blur loss ($\mathcal{B}_{b}$) and 2048 rays for the event loss ($\mathcal{B}_{e}$), and sample 64 coarse and 64 fine points along each ray. The number of motion samples $M$ per exposure is scene-dependent and typically ranges from 7 to 11 based on motion complexity. The pose refiner $f_{\phi}$ is implemented as an 8-layer MLP with 256-dimensional hidden units and ReLU activations, following [28], while we follow [4] for implementing the fine and coarse MLPs defining the NeRF. We use the Adam optimizer [18] to minimize the total loss $\mathcal{L}=\lambda_{b}\mathcal{L}_{\text{blur}}+\lambda_{e}\mathcal{L}_{\text{event}}+\lambda_{p}\mathcal{L}_{\text{prior}}$ where we set $\lambda_{b}=1.0$ and $\lambda_{e}=0.1$, while we decay $\lambda_{p}$ from an initial value of $0.1$. to zero by iteration 20,000 with a cosine scheduling to let the NeRF move beyond the model-based prior. We train for 30,000 iterations with a learning rate exponentially decaying from $5\cdot 10^{-3}$ to $5\cdot 10^{-6}$. The total runtime for training a single scene is around 3 hours on one NVIDIA A100 GPU.
@@ IV-D Experimental Validation
While precise trajectory recovery is not the primary objective of our method, these results demonstrate its ability to achieve accurate poses , exploiting dense 3D reconstruction as a way to jointly refine poses during training. Additional results are provided in Figure 4, where we qualitatively visualize the refined trajectories and report the changes in trajectory error with the distance traveled.
@@
The results are shown in Table V. Although using noisy UltimateSLAM poses [46] as initialization, our method maintains high performance across different exposure time and speed configurations, achieving 21.31 dB PSNR in the most challenging setting. In contrast, Ev-DeblurNeRF [4] using the same poses fails to produce competitive results in all settings due to the lack of a pose refinement stage. The performance gap between Ev-DeblurNeRF [4] and our method is consistently greater than 7 dB in PSNR. Crucially, even when provided with our refined poses, both Ev-DeblurNeRF [4] and E2NeRF [37] baselines consistently underperform relative to our approach. The first is limited to linearly interpolating poses at event timestamps, while the second only considers events during image exposures, thus underutilizing high-frequency supervision. Qualitative results illustrating these findings are in Figure 7.
@@
Adding the event loss on top of the blur loss yields large improvements (+20% PSNR, +21% LPIPS, +31% SSIM on average), confirming the strong benefit of event supervision. When using only the prior loss, compared to using only the blur loss, the performance gap increases significantly from +0.21dB at 1 m/s to +5.51dB at 2 m/s case, with an average gap of +2.53dB PSNR, demonstrating the usefulness of the pseudo ground-truth target obtained from the model-based deblurring process, particularly in high-speed scenarios. Similarly, after adding the event loss on top of the prior loss, the average performance increases considerably (+7% PSNR, +21% LPIPS and +17% SSIM), which again demonstrates the significance of event supervision. Combining the blur loss and the prior loss yields better results than using either loss alone. Nonetheless, configurations that include the event loss (e.g., blur+event or prior+event) consistently outperform the blur+prior combination. The highest performance is achieved when all three loss terms are used jointly. Finally, removing the eCRF module leads to noticeable performance degradation (-2% PSNR, -11% LPIPS, -3% SSIM), demonstrating the effectiveness of modeling the event camera response function.
@@
BoxStack USLAM EvDeblurNeRF [4] 13.34 0.58 0.30 12.08 0.64 0.25 13.58 0.66 0.30 13.00 0.63 0.28
@@
Equipment USLAM EvDeblurNeRF [4] 13.86 0.53 0.30 13.49 0.63 0.27 13.08 0.73 0.24 13.48 0.63 0.27
@@
PlayCorner USLAM EvDeblurNeRF [4] 13.30 0.70 0.31 11.66 0.78 0.20 11.92 0.76 0.26 12.29 0.75 0.26
@@
Models USLAM EvDeblurNeRF [4] 10.50 0.74 0.19 10.71 0.77 0.16 11.48 0.78 0.18 10.90 0.76 0.18
@@
ReadingCorner USLAM EvDeblurNeRF [4] 13.23 0.69 0.36 13.76 0.67 0.39 11.77 0.70 0.13 12.92 0.69 0.29
@@
Lounge USLAM EvDeblurNeRF [4] 12.26 0.68 0.17 12.65 0.69 0.18 11.74 0.69 0.13 12.22 0.69 0.16
@@
Average USLAM EvDeblurNeRF [4] 12.75 0.65 0.27 12.39 0.70 0.24 12.26 0.72 0.21 12.47 0.69 0.24
@@
TABLE VII: Quantitative comparison on trajectories of different lengths. All trajectories flew by real drone under the fastest speed profile. Scene 1: Models, Scene 2: ReadingCorner.
@@ V Discussion and Limitations
Extension to Gaussian Splatting.  Our focus in this work is accurate and stable reconstruction under the fast motion of aerial drones, where robustness to sparse input frames and generalization to unseen viewpoints are critical due to high motion speed and limited capture frequency. Recent analyses [15, 10] indicate that NeRFs are more stable to train and provide stronger geometric consistency and better generalization than 3D Gaussian Splatting with limited training views. Our choice of a NeRF backbone therefore prioritizes reconstruction fidelity and robustness under sparse training views. However, our framework is agnostic to the underlying representation of the scene. In scenarios where rendering speed is more desired, our method can be naturally extended to 3DGS. Compared to NeRF which represents the scene implicitly through a neural network, 3DGS models it explicitly as a set of Gaussian splats parameterized by position, covariance, and appearance. By rendering these splats differentiably with respect to the continuous-time camera trajectory analogous to our current formulation, the same image- and event-based supervision can be used to jointly optimize both the splat parameters and the camera poses.

### v2

@@ III-A Trajectory Representation and Pose Refinement
where $\gamma_{L_{t}}(\cdot)$ denotes a frequency-encoded representation of time, following the sinusoidal positional encoding introduced in NeRF [29]. The network $f_{\phi}$ maps this embedding to the residual pose $\Delta\mathbf{T}_{\text{refine}}(t)$, computed from the predicted translation vector $\mathsf{v}(t)\in\mathbb{R}^{3}$ and quaternion $\mathsf{q}(t)\in\mathbb{R}^{4}$.
@@ III-B Image Supervision via Motion Blur Rendering
Given a pixel $\mathbf{u}$ in a frame captured at time $t_{i}$, where $t_{i}$ denotes the center of the exposure interval for image $i$, we define $\mathbf{r}(\mathbf{u},t_{i})$ as the ray that originates from the camera center of the refined camera pose $\mathbf{T}_{\text{ref}}(t_{i})$ and passes through the pixel $\mathbf{u}$. We estimate the color of the blurred pixel $\mathbf{u}$ observed in $\mathbf{C}^{\text{blur}}_{i}$ by averaging the colors observed along the set of rays $\mathbf{r}(\mathbf{u},t_{i})$ as the camera moves during the exposure interval $t_{i}-\tau/2,\ t_{i}+\tau/2]$. This simulates the physical image formation process, where the observed pixel color corresponds to the integral of the latent sharp radiance accumulated over the exposure duration [27].
@@
where $w_{m}$ denotes the blending weight for the $m$-th timestamp and satisfies $\sum_{m=1}^{M}w_{m}=1$, and $g(\cdot)$ is a gamma correction function. Similar to [37], we use uniform weights to blend the rendered colors over time.
For brevity, we denote the rendered pixel in Eq. 5 as $\hat{\mathbf{C}}^{\text{blur}}_{\mathbf{r}}$, we finally compare the synthetic and observed blurry pixels over a batch $\mathcal{B}_{b}$ of pixel-timestamp pairs $(\mathbf{u}_{k},t_{k})$:
@@ IV-A Implementation Details
Training.  We implement our code using PyTorch. We train using a batch size of 1024 rays for the blur loss ($\mathcal{B}_{b}$) and 2048 rays for the event loss ($\mathcal{B}_{e}$), and sample 64 coarse and 64 fine points along each ray. The number of motion samples $M$ per exposure is scene-dependent and typically ranges from 7 to 11 based on motion complexity. The pose refiner $f_{\phi}$ is implemented as an 8-layer MLP with 256-dimensional hidden units and ReLU activations, following [28], while we follow [4] for implementing the fine and coarse MLPs defining the NeRF. We use the Adam optimizer [18] to minimize the total loss $\mathcal{L}=\lambda_{b}\mathcal{L}_{\text{blur}}+\lambda_{e}\mathcal{L}_{\text{event}}+\lambda_{p}\mathcal{L}_{\text{prior}}$ where we set $\lambda_{b}=1.0$ and $\lambda_{e}=0.1$, while we decay $\lambda_{p}$ from an initial value of $0.1$ to zero by iteration 20,000 with a cosine scheduling to let the NeRF move beyond the model-based prior. We train for 30,000 iterations with a learning rate exponentially decaying from $5\cdot 10^{-3}$ to $5\cdot 10^{-6}$. The total runtime for training a single scene is around 3 hours on one NVIDIA A100 GPU.
@@ IV-D Experimental Validation
While precise trajectory recovery is not the primary objective of our method, these results demonstrate its ability to achieve accurate poses by exploiting dense 3D reconstruction as a way to jointly refine poses during training. Additional results are provided in Figure 4, where we qualitatively visualize the refined trajectories and report the changes in trajectory error with the distance traveled.
@@
The results are shown in Table V. Despite using noisy UltimateSLAM poses [46] as initialization, our method maintains high performance across different exposure times and speed configurations, achieving 21.31 dB PSNR in the most challenging setting. In contrast, Ev-DeblurNeRF [4] using the same poses fails to produce competitive results in all settings due to the lack of a pose refinement stage. The performance gap between Ev-DeblurNeRF [4] and our method is consistently greater than 7 dB in PSNR. Crucially, even when provided with our refined poses, both Ev-DeblurNeRF [4] and E2NeRF [37] baselines consistently underperform relative to our approach. The first is limited to linearly interpolating poses at event timestamps, while the second only considers events during image exposures, thus underutilizing high-frequency supervision. Qualitative results illustrating these findings are in Figure 7.
@@
Adding the event loss on top of the blur loss yields large improvements (+20% PSNR, +21% LPIPS, +31% SSIM on average), confirming the strong benefit of event supervision. When using only the prior loss, compared to using only the blur loss, the performance gap increases significantly from +0.21dB at 1 m/s to +5.51dB in the 2 m/s case, with an average gap of +2.53dB PSNR, demonstrating the usefulness of the pseudo ground-truth target obtained from the model-based deblurring process, particularly in high-speed scenarios. Similarly, after adding the event loss on top of the prior loss, the average performance increases considerably (+7% PSNR, +21% LPIPS and +17% SSIM), which again demonstrates the significance of event supervision. Combining the blur loss and the prior loss yields better results than using either loss alone. Nonetheless, configurations that include the event loss (e.g., blur+event or prior+event) consistently outperform the blur+prior combination. The highest performance is achieved when all three loss terms are used jointly. Finally, removing the eCRF module leads to noticeable performance degradation (-2% PSNR, -11% LPIPS, -3% SSIM), demonstrating the effectiveness of modeling the event camera response function.
@@
BoxStack USLAM Ev-DeblurNeRF [4] 13.34 0.58 0.30 12.08 0.64 0.25 13.58 0.66 0.30 13.00 0.63 0.28
@@
Equipment USLAM Ev-DeblurNeRF [4] 13.86 0.53 0.30 13.49 0.63 0.27 13.08 0.73 0.24 13.48 0.63 0.27
@@
PlayCorner USLAM Ev-DeblurNeRF [4] 13.30 0.70 0.31 11.66 0.78 0.20 11.92 0.76 0.26 12.29 0.75 0.26
@@
Models USLAM Ev-DeblurNeRF [4] 10.50 0.74 0.19 10.71 0.77 0.16 11.48 0.78 0.18 10.90 0.76 0.18
@@
ReadingCorner USLAM Ev-DeblurNeRF [4] 13.23 0.69 0.36 13.76 0.67 0.39 11.77 0.70 0.13 12.92 0.69 0.29
@@
Lounge USLAM Ev-DeblurNeRF [4] 12.26 0.68 0.17 12.65 0.69 0.18 11.74 0.69 0.13 12.22 0.69 0.16
@@
Average USLAM Ev-DeblurNeRF [4] 12.75 0.65 0.27 12.39 0.70 0.24 12.26 0.72 0.21 12.47 0.69 0.24
@@
TABLE VII: Quantitative comparison on trajectories of different lengths. All trajectories were flown by a real drone under the fastest speed profile. Scene 1: Models, Scene 2: ReadingCorner.
@@ V Discussion and Limitations
Extension to Gaussian Splatting.  Our focus in this work is accurate and stable reconstruction under the fast motion of aerial drones, where robustness to sparse input frames and generalization to unseen viewpoints are critical due to high motion speed and limited capture frequency. Recent analyses [15, 10] indicate that NeRFs are more stable to train and provide stronger geometric consistency and better generalization than 3D Gaussian Splatting with limited training views. Our choice of a NeRF backbone therefore prioritizes reconstruction fidelity and robustness under sparse training views. However, our framework is agnostic to the underlying representation of the scene. In scenarios where rendering speed is more important, our method can be naturally extended to 3DGS. Compared to NeRF, which represents the scene implicitly through a neural network, 3DGS models it explicitly as a set of Gaussian splats parameterized by position, covariance, and appearance. By rendering these splats differentiably with respect to the continuous-time camera trajectory analogous to our current formulation, the same image- and event-based supervision can be used to jointly optimize both the splat parameters and the camera poses.